OpenStudy (anonymous):

Which expression is equivalent to solving the quotient of two over nine division symbol five over seven? Answer two over nine times seven over five nine over two times five over seven nine over two times seven over five two over nine times five over seven

OpenStudy (anonymous):

\[\frac{2}{9}\div\frac{5}{7}=\frac{2}{9}\times \frac{7}{5}\]

OpenStudy (anonymous):

in general \[\frac{a}{b}\div\frac{c}{d}=\frac{a}{b}\times \frac{d}{c}\]